13. A blimp is flying parallel to a road and is 2460 feet directly above it. The angles of depression of two parked cars on the road are 34.9° and 26.5°. TO the nearest foot, how far apart are the parked cars? (Note: The blimp is above a line between the parked cars.
1
Expert's answer
2020-03-04T18:03:25-0500



Let L denote a  blimp, A and B - cars. Then "LR = 2460"  feet , "\varphi_1 = 26.5°, \varphi_2 = 34.9°." The goal is to find distance "AB = AR + RB.\\"

From the triangle "ALR" : "AR = LR\cdot \tan(90°-\varphi_1) \approx 2460\cdot 2=4920\\"

From the triangle "BLR" : "RB = LR\cdot \tan(90°-\varphi_2) \approx 2460\cdot1.43\approx3526.3\\"

Finaly, "AB = AR + RB = 4920+3526.3 \approx 8446.3 \approx 8446.\\"

Answer: The parked cars are 8446 feet apart.